OpenStudy (ritoking):

What is the slope-intercept form of the linear equation 4x + 2y = 24?

OpenStudy (3mar):

It is in the general form \[y=mx+b\]

OpenStudy (rcp031):

Standard form is y=mx+c here m=slope=rise/run, c=constant

OpenStudy (mhchen):

Slope intercept form is y = mx + b So you have 4x + 2y = 24 We can do: 2y = 24 - 4x y = (24/2) - (4x/2) <--simplify this

OpenStudy (rcp031):

transpose it in y form